What must be the magnitude and direction of E that will balance the weight of

a.) an electron

b.) a proton

show your solution


1
Expert's answer
2022-03-21T12:44:22-0400

We know that

Electric force =weight

"qE=mg"

"E=\frac{mg}{q}"

(a)

For electron


"E=-\frac{9.1\times10^{-31}\times9.8}{1.6\times10^{-9}}=-5.57\times10^{-11}N/C\\|E|=5.57\times10^{-11}N/C"

Direction will be downward charge is negative

(B)

For proton


"E=\frac{1.67\times10^{-27}\times9.8}{1.6\times10^{-19}}=1.022\times10^{-7}N/C"

direction will be upward charge for positive